OpenStudy (anonymous):

integrate f(q)= (q -4)² - 12

OpenStudy (anonymous):

why not expand the brackets first then differentiate?

OpenStudy (anonymous):

expand the brackets first then integrate

OpenStudy (anonymous):

(a-b)^2 = a^2-2ab + b^2

OpenStudy (anonymous):

Nope no need to expand. \[\int\limits [(q-4)^2-12] dq\] \[\frac{1}{3}(q-4)^3-12q+C\]

OpenStudy (anonymous):

Anti-differentiate.

OpenStudy (anonymous):

Anti-differentiating is one of the fastest ways to integration unless there comes a time when anti-differentiation starts to not work, then you go to other methods (eg. Using partial fractions" to integrate).